OpenStudy (yanasidlinskiy):

Did I set this equation up correctly?? Let f(x) = -2x - 7 and g(x) = -4x + 3. Find (f o g)(-5) A. 3 B. 23 C. -9 D. -53 (-2x-7)*(-4x+3)(5) <----That's what I did.

OpenStudy (yanasidlinskiy):

@ikram002p

OpenStudy (ikram002p):

f o g)(-5) means f after g of -5 (f o g)(-5) =f (g(-5) ) first find g(-5) could you do that ?

OpenStudy (yanasidlinskiy):

I don't even know.

OpenStudy (ikram002p):

g(x) = -4x + 3 to find g(-5) replace x with -5

OpenStudy (yanasidlinskiy):

Ok. It's g(-5)=-4x+3

OpenStudy (yanasidlinskiy):

Let me solve it and tell you what I got.

OpenStudy (ikram002p):

well on the other side also u have x replace it with -5 g(-5)=-4(-5)+3

OpenStudy (yanasidlinskiy):

oh, wait. just Kidding. I don't know what to do next. I'm kind of stuck.

OpenStudy (yanasidlinskiy):

Ok. This is what I got. G-5=23

OpenStudy (yanasidlinskiy):

Would I add 5 to both sides?

OpenStudy (ikram002p):

no its enough u have g(-5) is 23 :)

OpenStudy (yanasidlinskiy):

Ok. Now what do I do?

OpenStudy (ikram002p):

ok nw lwts continue (f o g)(-5) =f (g(-5) ) = f(23 ) right ? find f(23 ) f(x)== -2x - 7

OpenStudy (yanasidlinskiy):

So..I would replace the x with 23 right?

OpenStudy (ikram002p):

sure :)

OpenStudy (yanasidlinskiy):

f(23)=-2(23)-7 f(23)=-46-7 f(23)=-53. Is that right?

OpenStudy (ikram002p):

that is correct \(\checkmark \)

OpenStudy (yanasidlinskiy):

Haha:D Ok. Is that all what we have to do? Or is there more to it?

OpenStudy (ikram002p):

no its all/;p

OpenStudy (yanasidlinskiy):

Hahha!!:D Ok. I thought it was like some kind of long problem;) Thanx sooooooooooo much for your time!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! \(\huge\cal\color{lime}{Thank~you!!!}\) \(\huge{\leftarrow}\)

OpenStudy (ikram002p):

np :)